# Generative Models in Decision Making: A Survey

### v2

@@ Abstract
In recent years, the exceptional performance of generative models in generative tasks has sparked significant interest in their integration into decision-making processes. Due to their ability to handle complex data distributions and their strong model capacity, generative models can be effectively incorporated into decision-making systems by generating trajectories that guide agents toward high-reward state-action regions or intermediate sub-goals. This paper presents a comprehensive review of the application of generative models in decision-making tasks. We classify seven fundamental types of generative models: energy-based models, generative adversarial networks, variational autoencoders, normalizing flows, diffusion models, generative flow networks, and autoregressive models. Regarding their applications, we categorize their functions into three main roles: controllers, modelers and optimizers, and discuss how each role contributes to decision-making. Furthermore, we examine the deployment of these models across five critical real-world decision-making scenarios. Finally, we summarize the strengths and limitations of current approaches and propose three key directions for advancing next-generation generative directive models: high-performance algorithms, large-scale generalized decision-making models, and self-evolving and adaptive models.
@@ 1 Introduction
Generative models have become a hot topic in both academia and industry, primarily due to their ability to generate large quantities of synthetic data with high quality and diversity. From early systems like DALL-E [1] (for image generation) and GPT-3 [2] (for text generation) to more recent advancements such as DALL-E3 [3], ChatGPT and GPT-4 [4], generative models have rapidly advanced in both quality and scale of their outputs.
@@
Content generation aims to create coherent material that mimics training examples, while decision-making focuses on producing action sequences for optimal outcomes. Unlike content generation, decision-making involves complex, dynamic environments and long-term decisions. Thus, despite generative models’ success in content generation, applying them to decision-making poses challenges. These challenges include: 1) how to learn policies through interaction with the environment, rather than simply mimicking expert behavior. 2) how to generate new policies based on learned behaviors, transitioning from policy learning to policy generation. 3) how to establish a robust fundamental decision generating model that can adapt to various environments with minimal tuning efforts. 4) how to build multi-step reasoning and long-term evolution capabilities of strategies. These challenges emphasize the need for generative models to go beyond mere data generation.
In practice, decision-making is often referred to as sequential decision making, where a decision maker makes a series of observations over time, and each decision influencs subsequent choices. The goal is to identify a policy that optimizes expected rewards or minimizes costs across sequential actions. Classical algorithms such as Dynamic Programming (DP) and Reinforcement Learning (RL) are widely used to solve problems modeled as Markov Decision Processes (MDPs). These methods optimize decision-making by updating policies based on observed rewards and state transitions rather than generating new ones. Despite their many successful applications, these traditional approaches often rely on trial-and-error or predefined states and transitions, limiting exploration and potentially missing better solutions. Moreover, they require substantial computation and optimization, which can be impractical for high-dimensional or large-scale problems. Traditional methods also need significant reconfiguration or retraining to new environment, reducing flexibility.
On the other hand, generative models are designed to model data distributions, rather than simply fitting labels. Once trained, they can generate new samples that resemble the original data, enabling the exploration of diverse scenarios and outcomes. This ability enables the discovery of novel strategies that may not be immediately evident with traditional methods. In complex or poorly labeled data scenarios, generative models offer a richer understanding of possible decision paths, sometimes leading to strategies that better align with high rewards or desired goals. However, traditional methods like optimization or reinforcement learning remain effective in simpler, well-defined environments where the decision space is clearer, and goals are more straightforward. The choice between these approaches depends on the task complexity and the environment characteristics.
@@
Recognizing these advantages, recent years have seen substantial research efforts aimed at developing new generative models and applying them to decision making. Fig. 1 illustrates the research trends in generative models and their applications in decision making, further emphasizing the significance of these methods in addressing such challenges. However, there is a lack of comprehensive reviews that summarize past work and pave the way for new research avenues. This gap motivates us to provide this review paper.
@@
The survey highlights three major contributions: 1) A comprehensive taxonomy is proposed for classifying current generative decision-making methods. We identify seven types of generative models used in decision-making and categorize their functions into three key roles: controller, modeler, and optimizer. 2) We review the diverse practical uses of generative models in decision-making, focusing on robot control, structural generation, games, autonomous driving and optimization tasks. 3) Finally, we summarize the advantages and limitations of existing work and discuss future perspectives for developing high-performance generative models in decision-making tasks.
@@
The organization of the rest of this survey is as follows (refer to Fig. 2 for a general outline): Section 2 serves as a preliminary by introducing sequential decision-making formulation and provides the basics of all the examined methods. Specifically, we offer a detailed introduction to seven types of generative models and compare their performance with traditional approaches. Section 3 presents the proposed taxonomy for categorizing generative decision-making methods. In Section 4, we review and analyze existing literature according to the introduced taxonomy. Section 5 showcases practical applications of generative models in decision-making. Finally, Section 6 discusses the future directions of generative models in decision-making, and we conclude the paper in Section 7 with an overall summary.
@@
Figure 1: Research trends in generative models and their applications in decision-making (2000-2024). The bars represent the rough average annual number of papers, sourced from Google Scholar. The search included titles associated with seven classic types of generative models and their applications in five real-world scenarios we detailed in Fig. 20.
@@ 2.1 Sequential Decision Making
Sequential decision-making involves making step-by-step decisions, where each choice depends on previous outcomes. At each step, the agent typically observes the current state, selects an optimal action based on its policy, and the environment updates its state while providing new rewards. The ultimate goal is to maximize the accumulated rewards.
The Sequential decision process is categorized according to the number of agents and the observation to the environment. In this paper, we only focus on the single agent. If the agent is able to fully observe the environment, it is considered a Markov Decision Process (MDP) [5]. An MDP is a framework for modeling discrete-time decision-making process, where outcomes depend on both random factors and the agent’s decisions [5]. Formally, an MDP is defined as a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{R},P,\rho_{0},\gamma,H)$ to represent MDP, where $\mathcal{S}$ is the state space, and $s\in\mathcal{S}$ contains all the information perceived by the agent from the environment. $\mathcal{A}$ represents the action space, encompassing all possible actions $a$ that the agent can execute when interacting with the environment. $r\in\mathcal{R}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ denotes the reward function, which corresponds to the transition pair $(s,a,s^{\prime})$. $P:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ refers to the transition function. When the action $a$ is applied to the state $s$, $P(s^{\prime}|s,a)$ generates a distribution over states. $\rho_{0}\in\Delta(\mathcal{S})$ represents the initial state distribution. The discount factor $\gamma\in[0,1]$ quantifies the long-term value of the current action, while $H$ denotes the horizon.
@@ 2.2 Related Methods
We introduce several traditional methods for solving sequential decision making problems in this section.
2.2.1 Imitation Learning
Imitation learning aims to optimize a policy $\pi_{\theta}(s)$ such that the action distribution is close to that of an expert policy $\pi^{*}$ for any state $s$, the objective function is given by
$\theta=\arg\min_{\theta}\mathbb{E}_{s\sim p(s|\pi_{\theta})}L\left(\pi^{*}(s)-\pi_{\theta}(s)\right),$ (3)
where $p(s|\pi_{\theta})$ denotes the state distribution induced by $\pi_{\theta}$, and $L$ is a metric function.
2.2.2 Search based Methods
Policy search focuses on exploring the policy space directly, rather than computing the value function explicitly. For a given policy $\pi$, the expected discounted return is given by
$\mathcal{U}(\pi)=\sum_{s}\rho_{0}(s)\mathcal{U}^{\pi}(s).$ (4)
With large state space, $\mathcal{U}(\pi)$ could be approximated by sampling trajectories consists of $(\mathcal{S},\mathcal{A},\mathcal{R})$ pairs and $\mathcal{U}(\pi)$ can be reformulated as [7]:
[…]
where $R(\tau)$ denotes the discounted return related to $\tau$. Monte Carlo policy evaluation entails estimating the expected utility of the policy $\pi$ by performing numerous rollouts starting from $s_{0}\sim\rho_{0}(s)$. Local search starts with a feasible solution and incrementally moves to the neighbor with better utility over the search space until convergence occurs, e.g. Hooke-Jeeves method [8]. Simulated annealing [9] allows for occasional moves to feasible solutions with worse utility to approximate the global optimum. Genetic algorithms evaluate different simulations concurrently based on the objective, recombine them and guide the population toward a global optimum.
2.2.3 Planning & Optimization
Planning is an optimization approach that relies on a predefined model to guide the search process. Let $U(s)$ denote the action space for each state $s$, which represents the set of all actions that could be applied from $s$. For distinct $s,s^{\prime}\in\mathcal{S}$, $U(s)$ and $U(s^{\prime})$ are not necessarily disjoint. As part of the planning problem, a set $\mathcal{S}_{G}\subset\mathcal{S}$ of goal states and the initial state $s_{I}$ are defined.
Let $\pi_{K}$ denote a $K$-step plan, which is a sequence of $(a_{1},a_{2},\cdots,a_{K})$ of $K$ actions. Let $F$ denote the final stage where $F=K+1$. The cost functional is defined as
$L(\pi_{K})=\sum_{k=1}^{N}l(s_{k},a_{k})+l_{F}(s_{F}),$ (6)
where $l(s_{k},a_{k})$ is the cost term yielding a real value for every $s_{k}\in\mathcal{S}$ and $a_{k}\in U(s_{k})$. The final term is defined $l_{F}(s_{F})=0$ if $s_{f}\in\mathcal{S}_{G}$ and $l_{F}(s_{F})=\infty$ otherwise.
@@ 2.2.4 Reinforcement Learning
By introducing KL regularization and clipped surrogate objective, True Region Policy Optimization (TRPO) [16] and Proximal Policy Optimization (PPO) [17] constrain policy size updated at each iteration to stable the training process. These methods and their variants have achieved great success in areas like robot control [18], games [19] and generative model training [20].
Actor-Critic Methods. Taking advantage of both policy gradient and value-based approaches, Actor-critic algorithms use a parameterized policy and value function. And the value function can provide a better estimation of $\hat{A}(s_{t},a_{t})$ for policy gradients, with calculating the average return over the sampled trajectories.
Unlike Q-learning, the actor-critic method seeks to optimize the Q-function associated with the current policy $\pi_{\theta}$ rather than learning the optimal Q-function directly. Many variants are proposed based on this principle, such as on-policy and off-policy algorithms [21]. Deep Deterministic Policy Gradient (DDPG) introduces a deterministic policy under the Actor-Critic architecture, which can solve continuous control problems [21]. Soft actor-critic introduce the entropy regularization into the policy optimization process, so that the agent can make better exploration [22].
Model-based RL. For model-based RL, the first step is to learn the environment model $P(\cdot|s,a)$ based on historical trajectories. According to the probabilistic transition model, a common approach is to reduce the KL divergence between the learned model $\hat{P}(\cdot|s,a)$ and the true dynamics $P(\cdot|s,a)$ as [23]:
@@
where $\theta$ denotes the learned model’s parameters. The model learning process can be transformed as a supervised learning task, which can be solved effectively by some supervised learning technique. To reduce the error of model learning, several scholars propose the Lipschitz continuity constraints[24], distribution matching[25], and robust model learning methods[26] for further precise estimation.
When the model is ready, the agent develops plans to improve its strategy to interact with the modeled world. Traditional methods for integrating planning into MBRL mainly include model predictive control, Monte Carlo tree search and Dyna-style methods [27]. Dyna-style methods utilize the learned model to generate more experience and then perform RL on the model-augmented dataset, which has been the mainstream method [28]. Similarly, these methods employ value function estimation and policy gradient to improve the policy.
@@ 2.3 Generative Models
Generative Model  [29, 30, 31, 32, 33, 34, 35, 36] is an important subdivision of artificial intelligence that employs techniques and models specifically designed for sampling unseen data from the underlying distribution of the existing dataset.
In this section, we explore generative models through three crucial dimensions: sample quality, sample diversity, and computational efficiency [37, 38]. These dimensions are essential for understanding how generative models perform in decision-making, as they directly impact the accuracy, robustness, and practical applicability of the generated outputs. Sample quality measures how well the generated samples align with the real data distribution, reflecting the realism and reliability of the outputs. High sample quality ensures that decisions based on the generated data are both accurate and trustworthy in real-world contexts. Sample diversity, on the other hand, evaluates how well the model can generate a broad range of distinct samples, capturing the full spectrum of potential outcomes. This is vital in decision-making tasks, where diverse options are necessary to adapt to varying conditions and avoid narrow, overfitted solutions. Finally, computational efficiency refers to the computational resources required for training and inference, making it a key consideration for the deployment of generative models in real-world applications. In contexts such as autonomous driving or robotic control, where decisions must be made in real time and under resource constraints, computational efficiency is critical to ensure that the model remains practical and scalable. Together, these dimensions provide a comprehensive framework for evaluating generative models, guiding their application in complex, resource-sensitive decision-making environments.
We will examine each of the seven generative models through these three crucial dimensions, identifying their strengths and weaknesses. Generative models face challenges in balancing sample quality, sample diversity, and computational efficiency. For example, Diffusion Models and Normalizing Flows offer strong sample diversity and stability but require high computational resources, limiting their suitability for real-time decision-making applications [39, 40, 41]. In contrast, models like VAEs and GANs provide faster training and better efficiency but may struggle with maintaining sample diversity, potentially leading to overly similar or overfitted outputs [42, 43, 34, 44]. Based on the references [45, 46, 43, 47, 48, 49, 50, 51, 52, 53], we compare the performance of these seven generative models across sample quality, diversity, and efficiency, as shown in Fig. 3.
@@
Variational Autoencoders (VAEs). VAEs  [43], known for their stable learning process and efficient computation, attract significant attention in recent years. VAEs [68] are probabilistic models that represent the input data in a compressed form, referred to as the latent space. New samples are generated by sampling from this latent space and then decoding them back into the original data domain. The goal of a VAE, presented in (15), aims to maximize the evidence lower bound (ELBO):
@@
VAEs are deployed to generate new images or videos[69], and augment images, such as image inpainting [70] and image colorization [71]. VAEs are used to compress data [72, 73] or provide a good way to detect anomalies or outliers in data [74]. However, VAEs may suffer from low generation quality and the independence assumption between the latent variables [72]. To address these limitations, several extensions of VAEs have been proposed, e.g., conditional VAE [75] and Gaussian mixture VAE[76].
Researchers propose to combine VAEs with other techniques, including adversarial training [77, 78] and normalizing flows [79] for improving disentanglement. Besides,  [80] and  [81] study the continual learning of VAE representations and  [82] trains a response function of the hyperparameter $\beta$ in $\beta-VAE$ to trade off the reconstruction error and the KL divergence.
Normalizing Flows (NFs). Flow-based models [83] learn to map a simple random variable, such as a Gaussian distribution, to a more complex distribution capable of representing data. Normalizing Flow  [84, 85] applys a specific type of invertible transformation to the simple distribution to preserve the tractability of the density function, as shown below:
@@
A primary strength of normalizing flows is their precision in capturing complex probability. Normalizing flows can also be used to generate new samples and to perform density estimation, among other tasks. Despite their success, normalizing flows still face some challenges, such as the restricted network structure and the rendered limited expressiveness. Nevertheless, Normalizing flows remain a dynamic field of study, with continuous advancements being made.
From limited pre-neural attempts [86, 87, 88], specific flow transformations [89] are proposed to reduce the computational cost, such as NICE [90], Real-NVP [91], and Glow [41]. Researchers also use autoregressive flows [92, 93, 94] to accelerate the sampling process. Many Normalizing Flow variants also leverage the techniques of Neural ODE [95, 96, 97] and optimal transport [98, 99].
Diffusion Models (DMs). Diffusion models, introduced by Sohl-Dickstein et al. in 2015 [36], learn to generate samples through a diffusion process. This process involves gradually introducing noise to the data, followed by the reversal of this noise addition during sampling to recover the original data (as illustrated in Fig. 4). The key strength of diffusion models lies in their ability to generate high-quality, realistic, and diverse outputs, surpassing earlier generative models such as GANs in various image synthesis tasks [100]. This capability has sparked significant interest in their application to decision-making scenarios, particularly in high-dimensional environments.
@@
The reverse process can be computationally expensive, as it involves many time steps, each requiring careful sampling to accurately reconstruct the data. Despite this challenge, diffusion models can generate high-quality samples, particularly in fields such as image generation, where they have outperformed GANs in terms of visual fidelity and diversity.
In the context of decision-making, diffusion models offer several advantages. They can model complex, high-dimensional decision spaces and generate solutions that explore a wide range of potential outcomes, making them suitable for applications in reinforcement learning (RL), robotics, and automated planning.
For instance, in robotic control, diffusion models can generate motion sequences that not only mimic the desired behavior but also explore alternative strategies that could lead to better performance in uncertain or dynamic environments. Similarly, in autonomous driving, they can generate realistic trajectories and responses to dynamic driving scenarios, taking into account the uncertainty of the environment.
Moreover, diffusion models can facilitate the exploration of decision spaces in complex environments where traditional models might struggle to account for the full range of possibilities. This capability is particularly useful in multi-objective decision making, where a model needs to balance trade-offs between conflicting goals.
Diffusion models demonstrate remarkable potential for producing high-quality, realistic, and novel data, and they continue to be an important tool in the field of generative modeling. Nevertheless, diffusion models have a slow sampling process, which will limit its scalability to high-dimensional cases. Consistency models [102] are a potential solution that support fast one-step generation.
@@
Autoregressive Models (AMs). Autoregressive models are commonly used in natural language processing and generative modeling tasks. They are a key component of sequence-to-sequence models, such as the Transformer model[109], and are responsible for generating output sequences based on an input sequence or a given context.
Autoregressive models operate by predicting each output element sequentially conditioned on the preceding ones. This means that the generation process is sequential, with the model generating the elements in a left-to-right fashion. At each step, the model takes into account the previously generated elements and the context information to make predictions.
A popular method for autoregressive generation is using an autoregressive language model, such as GPT [110], which has proven successful in NLP tasks like text generation, machine translation, and dialogue systems.
In autoregressive models, the input to each step is typically a combination of the previously generated elements and a context vector that captures the overall context or representation of the input sequence. The context vector can be obtained from an encoder network that processes the input sequence and produces a fixed-size representation, which is then used as input to the decoder.
During training, autoregressive models are typically optimized using Maximum Likelihood Estimation (MLE) [111]. The model aims to optimize the likelihood of generating target output sequence given the input sequence or context. This involves calculating the probability of each target element based on the preceding elements and updating parameters to improve its predictions [112].
Overall, autoregressive models are a powerful tool for sequence generation tasks, particularly in natural language processing [113]. They have contributed to significant advancements in generative modeling and continue to be an active area of research in the field.

### v3

@@ Abstract
In recent years, the exceptional performance of generative models in generative tasks has sparked great interest in their integration into decision-making processes. Due to their ability to handle complex data distributions and their strong model capacity, generative models can be effectively incorporated into decision-making systems by generating trajectories that guide agents toward high-reward state-action regions or intermediate sub-goals. This survey offers an in-depth overview of how generative models are applied to decision-making tasks. We categorize seven core types of generative models: energy-based models, generative adversarial networks, variational autoencoders, normalizing flows, diffusion models, generative flow networks, and autoregressive models. In terms of their applications, we divide their functions into three main roles: controllers, modelers and optimizers, and discuss how each role contributes to decision-making. Additionally, we explore how these models are applied in five key real-world decision-making situations. In conclusion, we highlight the strengths and limitations of existing approaches and suggest three key directions for advancing next-generation generative directive models: enhanced performance algorithms, large-scale generalized decision-making models, and self-evolving adaptive models. A curated list of papers categorized by our taxonomy is available at https://github.com/xyshao23/Awesome-Generative-Models-for-Decision-Making-Taxonomy.
@@ 1 Introduction
Generative models have seen rapid advancements, from DALL-E [1] (for image generation) and GPT-3 [2] (for text generation) to recent iterations such as DALL-E 3 [3], ChatGPT, and GPT-4 [4]. Despite their exceptional performance in content generation, applying them to decision-making poses several challenges, including: 1) How to learn policies through interaction with the environment, rather than simply mimicking expert behavior? 2) How to generate new policies based on learned behaviors, transitioning from policy learning to policy generation? 3) How to develop a robust decision-making model that can adapt to various environments with minimal tuning efforts? 4) How to build multi-step reasoning and long-term evolution capabilities of strategies? These challenges highlight the necessity for generative models to go beyond simple data generation.
In practice, decision-making is often framed as sequential decision-making, where each decision influences future actions. Classical methods like Dynamic Programming (DP) and Reinforcement Learning (RL) optimize policies based on observed rewards and state transitions in Markov Decision Processes (MDPs). However, they often rely on trial-and-error or predefined states, which limits exploration. Additionally, they can be computationally expensive and require retraining when applied to new environments. In contrast, generative models aim to capture data distributions and generate new samples, which enhances the exploration of diverse decision paths. This makes them particularly valuable in complex environments where traditional methods may fall short. However, in simpler, well-defined environments, classical approaches remain effective.
@@
Building on the advantages highlighted above, recent years have seen substantial research efforts focused on advancing generative models and applying them to decision-making. Fig. 1 illustrates the research trends in generative models and their applications in decision-making, further underscoring the significance of these methods in addressing such challenges. However, a thorough review summarizing past work and identifying new research directions is still lacking. This gap motivates us to provide this survey.
@@
The survey presents three major contributions: 1) We propose a comprehensive and detailed taxonomy for classifying generative decision-making methods, identifying seven types of models and categorizing their roles into controller, modeler, and optimizer. 2) We review the practical and diverse uses of generative models in decision-making, focusing on robot control, structural generation, games, autonomous driving, and optimization tasks. 3) We summarize the key advantages and limitations of existing work and discuss future perspectives for developing high-performance generative models in decision-making.
@@
The organization of the rest of this survey is as follows (refer to Fig. 2 for a general outline): Section 2 serves as a preliminary by introducing sequential decision-making formulation and provides the basics. Specifically, we offer a detailed introduction to seven types of generative models and compare their performances with traditional approaches. Section 3 presents the proposed taxonomy for categorizing generative decision-making methods. In Section 4, we review and analyze existing literature according to the introduced taxonomy. Section 5 showcases practical applications of generative models in decision-making. Finally, Section 6 discusses the future directions of generative models in decision-making, and we conclude in Section 7.
@@
Figure 1: Research trends in generative models and their applications in decision-making (2000-2024). The bars represent the rough average annual number of papers, sourced from google scholar. The search includes titles associated with seven classic types of generative models and their applications in five real-world scenarios we details in Fig. 15.
@@ 2.1 Sequential Decision Making
Sequential decision making involves step-by-step decisions where each choice depends on prior outcomes. The agent observes the current state, selects an action based on its policy, and the environment updates the state while providing rewards. The goal is to maximize accumulated rewards.
If the agent can fully observe the environment, it is modeled as a Markov Decision Process (MDP) [5], which is defined as a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{R},P,\rho_{0},\gamma,H)$. Here, $\mathcal{S}$ is the state space, and $s\in\mathcal{S}$ contains all the information perceived by the agent from the environment. $\mathcal{A}$ represents the action space, encompassing all possible actions $a$ that the agent can execute when interacting with the environment. $r\in\mathcal{R}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ denotes the reward function, which corresponds to the transition pair $(s,a,s^{\prime})$. $P:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ refers to the transition function. When the action $a$ is applied to the state $s$, $P(s^{\prime}|s,a)$ generates a distribution over states. $\rho_{0}\in\Delta(\mathcal{S})$ represents the initial state distribution. The discount factor $\gamma\in[0,1]$ quantifies the long-term value of the current action, while $H$ denotes the horizon.
@@ 2.2 Related Methods
In this section, we introduce several traditional methods for solving sequential decision-making problems.
2.2.1 Dynamic Programming
Dynamic Programming (DP) is a method for solving sequential decision-making problems by breaking them into simpler subproblems. It is effective for problems with a known model and optimal substructure, such as Markov Decision Processes (MDPs). The goal in DP is to find an optimal policy $\pi^{*}(s)$ that maximizes the expected cumulative reward, achieved by solving the Bellman equations.
The Bellman equation for the value function $V(s)$ of a state $s$ is given by:
$V(s)=\max_{a}\left(R(s,a)+\gamma\sum_{s^{\prime}}P(s^{\prime}|s,a)V(s^{\prime})\right),$ (3)
where $R(s,a)$ is the immediate reward for taking action $a$ in state $s$, $\gamma$ is the discount factor, and $P(s^{\prime}|s,a)$ is the transition probability to the next state $s^{\prime}$.
Alternatively, for the action-value function $Q(s,a)$, the Bellman equation is:
$Q(s,a)=R(s,a)+\gamma\sum_{s^{\prime}}P(s^{\prime}|s,a)\max_{a^{\prime}}Q(s^{\prime},a^{\prime}),$ (4)
where $Q(s,a)$ represents the expected return from taking action $a$ in state $s$.
@@ 2.2.2 Reinforcement Learning
True Region Policy Optimization (TRPO) [13] and Proximal Policy Optimization (PPO) [14] stabilize training by introducing KL regularization and a clipped surrogate objective to limit policy updates at each iteration. These methods have been successful in areas such as robot control [15], games [16], and generative model training [17].
Actor-Critic Methods. Actor-critic methods combine policy gradients and value-based methods by using a parameterized policy and value function. The value function helps better estimate $\hat{A}(s_{t},a_{t})$ for policy gradients.
Unlike Q-learning, actor-critic methods focuses on optimizing the Q-function for the current policy $\pi_{\theta}$ rather than learning the optimal Q-function. Variants include on-policy and off-policy algorithms [18]. Deep Deterministic Policy Gradient (DDPG) introduces a deterministic policy for continuous control problems [18], while Soft Actor-Critic introduces entropy regularization to enhance exploration [19].
Model-based RL. For model-based RL, the first step is to learn the environment model $P(\cdot|s,a)$ based on historical trajectories. A common approach is to minimize the KL divergence between the learned model $\hat{P}(\cdot|s,a)$ and the true dynamics $P(\cdot|s,a)$ as [20]:
@@
where $\theta$ denotes the learned model’s parameters.
When the model is ready, the agent develops plans to improve its strategy to interact with the modeled world. Traditional methods for integrating planning into MBRL mainly include model predictive control, Monte Carlo tree search and Dyna-style methods [21]. Dyna-style methods utilize the learned model to generate more experience and perform RL on the model-augmented dataset, which has been the main method [22]. These methods also use value function estimation and policy gradients to refine the policy.
@@ 2.3 Generative Models
Generative models  [23, 24, 25, 26, 27, 28, 29, 30] are an crucial subfield of artificial intelligence, aimed at generating unseen data samples based on the underlying distribution of existing datasets. In this section, we explore generative models along three dimensions: sample quality, diversity, and computational efficiency [31, 32]. These dimensions are crucial for understanding generative model performance in decision-making, as they directly affect the accuracy, robustness, and applicability of generated outputs.
Balancing sample quality, diversity, and computational efficiency presents a challenge for generative models. For example, Diffusion Models and Normalizing Flows offer strong sample diversity and stability but require high computational resources, limiting their suitability for real-time decision-making applications [33, 34, 35]. In contrast, models like VAEs and GANs provide faster training and better efficiency but may struggle with maintaining sample diversity, potentially leading to overly similar or overfitted outputs [36, 37, 28, 38]. Drawing from the studies [39, 40, 37, 41, 42, 43, 44, 45, 46, 47], we roughly compare the performance of these seven generative models across sample quality, diversity, and efficiency, as shown in Fig. 3.
@@
Variational Autoencoders (VAEs). VAEs [37] are probabilistic models that learn to represent input data in a compressed latent space, from which new samples can be generated. The objective of VAEs is to maximize the evidence lower bound (ELBO):
@@
VAEs are used for generating and augmenting new data (e.g., images, videos) [62, 63, 64], data compression, and anomaly detection [65, 66]. However, they may suffer from low generation quality and the independence assumption between latent variables [65]. Extensions like conditional VAEs [67] and Gaussian mixture VAEs [68] have been proposed to address these issues. VAEs are also combined with techniques such as adversarial training [69, 70] and normalizing flowss [71] for improved performance, and research on continual learning and hyperparameter tuning (e.g., in $\beta$-VAE) [72] is ongoing.
Normalizing Flows (NFs). Flow-based models [73] map simple distributions, such as a Gaussian distribution, to a more complex distribution capable of representing data. Normalizing Flows [74, 75] apply invertible transformations to maintain a tractable density function, as in (2.3):
@@
Normalizing Flows excel in capturing complex probabilities and are used for tasks like sample generation and density estimation, though challenges such as restricted network structure and the rendered limited expressiveness remain. The field continues to evolve with advancements [76, 77, 78, 79], including methods like NICE [80], Real-NVP [81], Glow [35], autoregressive flows [82, 83, 84], and Neural ODEs [85, 86, 87] to enhance efficiency and expressiveness.
Diffusion Models (DMs). Diffusion models, introduced by Sohl-Dickstein et al. in 2015 [30], generate samples by progressively adding noise to data and then recovering the original data through a reverse diffusion process during sampling (as shown in Fig. 4).
@@
The reverse process, which involves multiple time steps, can be computationally expensive. However, it enables the generation of high-quality samples and has shown to outperform GANs in visual fidelity and diversity, particularly in image generation tasks [89].
Diffusion models are also well-suited for complex decision-making tasks, particularly in high-dimensional spaces. They can generate diverse solutions, making them ideal for applications in reinforcement learning (RL), robotics, and automated planning. For instance, in robotic control, diffusion models can generate motion sequences that explore alternative strategies for better performance in dynamic environments. Similarly, in autonomous driving, they can generate realistic trajectories while accounting for environmental uncertainties.
However, the slow sampling process limits scalability for high-dimensional tasks. Consistency models [90] offer a potential solution with fast, one-step generation.
@@
Autoregressive Models (AMs). Autoregressive models are widely used in natural language processing and generative modeling. They are central to sequence-to-sequence models, like the Transformer [97], which generate output sequences based on input or context. These models predict each output element sequentially, conditioned on the preceding ones, as shown in (17):
$p(y_{1},y_{2},\dots,y_{T})=\prod_{t=1}^{T}p(y_{t}\mid y_{1},y_{2},\dots,y_{t-1}).$ (17)
At each time step $t$, the model predicts the next element $y_{t}$ based on the context provided by the preceding elements.
A popular method for autoregressive generation is using autoregressive language models, such as GPT [98], which have proven successful in NLP tasks like text generation, machine translation, and dialogue systems.
In AMs, the input at each step combines the previous elements and a context vector, typically obtained from an encoder that processes the input sequence and generates a fixed-size representation for the decoder.
During training, AMs are optimized using Maximum Likelihood Estimation (MLE) [99], aiming to maximize the likelihood of generating the target sequence given the input or context. This involves calculating the probability of each target element based on preceding elements and updating model parameters to improve predictions [100].
Overall, AMs are a powerful tool for sequence generation tasks, particularly in NLP [101], significantly advancing generative modeling and remaining an active research area.